OpenStudy (spring98):

Which of the following would be equivalent to 3^2 • 3^5? 3^10/3^3 3^10/3^7 3^1*3^7 3^0

Nnesha (nnesha):

when we divide same bases we should `subtract` their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\] and when we `multiply` same bases we should `add` ther exponents \[\huge\rm x^m \times x^n=x^{m +n}\]

OpenStudy (spring98):

so would it be the third one?

OpenStudy (spring98):

from the answers?

Nnesha (nnesha):

3rd one is \[\huge\rm 3^1 \times 3^7\] ???

OpenStudy (spring98):

i know but i need the answer that is eqivelent to 3^2*3^5!

Nnesha (nnesha):

i said when we multiply same bases we should add their exponents so 3^2 times 3^5= ???

OpenStudy (spring98):

im not understanding it

Nnesha (nnesha):

here is an example \[\huge\rm 2^\color{ReD}{5} \times 2^\color{blue}{4} = 2^{\color{red}{5} +\color{blue}{4}} =2^9\]

Nnesha (nnesha):

so \[\huge\rm 3^\color{ReD}{2} \times 3^\color{blue}{5} = 3^{???}\]

OpenStudy (spring98):

3^7

Nnesha (nnesha):

yes right! now look at the options which option is equal to 3^7

OpenStudy (spring98):

the third option right?

Nnesha (nnesha):

here are some examples anything to the zero power is just equal to one \[\huge\rm (x)^0= 1~~~~(anything)^0=1\] and like i said when we divide same bases we should subtract their exponents example is \[\huge\rm \frac{ 2^5 }{ 2^3 }=2^{5-3}\]

Nnesha (nnesha):

\[3^1 \times 3^7=?\]would you add those exponents or subtract ?

OpenStudy (spring98):

subtract

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha when we divide same bases we should `subtract` their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\] and when we `multiply` same bases we should `add` ther exponents \[\huge\rm x^m \times x^n=x^{m +n}\] \(\color{blue}{\text{End of Quote}}\) read this

OpenStudy (spring98):

ok we add

Nnesha (nnesha):

right so \[\huge\rm 3^1 \times 3^7= 3^{1+7} \] is this equal to 3^7 ??

OpenStudy (spring98):

3^8

Nnesha (nnesha):

so that's not equal to 3^7 that's why 3rs one isn't correct

Nnesha (nnesha):

3rd*

Nnesha (nnesha):

try other options

OpenStudy (spring98):

o it would be the first one because you subtract it and it would equal 3^7 right?

Nnesha (nnesha):

:=) nice!!!

OpenStudy (spring98):

thanks @Nnesha you helped me alot!!! :)

Nnesha (nnesha):

my pleasure :=)